OpenStudy (anonymous):

How do you solve for x(t): dx/dt = ct + kx where c and x are constants? I'm not sure where to move from here because i have both t and x on the RHS please

OpenStudy (unklerhaukus):

i think you mean \(c\) and \(k\) are the constants\[\frac{\text dx}{\text dt} = ct + kx\]

OpenStudy (anonymous):

Yes, typo there, c and k are constants

OpenStudy (anonymous):

Is is allowed to leave x on the right side while you have dx on left?

OpenStudy (dumbcow):

isn't this a situation where you need to find the integrating factor \[\rightarrow e^{\int\limits_{}^{}k dt}\]

OpenStudy (dumbcow):

** sorry that should be a negative k

OpenStudy (unklerhaukus):

\[\frac{\text dx}{\text dt} - kx= ct \] integrating factor is \(R(x)=e^{\int-k\text dx}\)

OpenStudy (dumbcow):

here is a good explanation....scroll down to examples http://tutorial.math.lamar.edu/Classes/DE/Linear.aspx

OpenStudy (anonymous):

Thanks a lot for helping me, had no idea what was hapenning

OpenStudy (unklerhaukus):

\[R(x)=e^{-k\int\text dx}=e^{-kx}\]

OpenStudy (unklerhaukus):

\[\frac{\text dx}{\text dt} - kx= ct\] \[\frac{\text d}{\text dt}\left(xe^{-kx}\right) = cte^{-kx}\] \[xe^{-kx}=c\int te^{-kx}\text dt\]

OpenStudy (unklerhaukus):

\[xe^{-kx}=c\frac{ t^2}{2}e^{-kx}+C\]\[x=\frac{ct^2}{2}+Ce^{kx}\]